# Xingwang Li, , and Arumugam Nallanathan

### v1

@@ Abstract
The advent of Rydberg atomic quantum receivers (RAQRs) offers a new solution for the evolution of wireless transceiver architecture, promising unprecedented sensitivity and immunity to thermal noise. However, RAQRs introduce a unique non-linear signal model based on biased phase retrieval, which complicates fundamental channel estimation tasks. Traditional iterative algorithms often struggle in low signal-to-noise regimes and fail to capture complex and non-ideal system characteristics. To address this, we propose a novel model-driven deep learning framework for channel estimation in RAQRs. Specifically, we propose a Transformer-based unrolling architecture, termed URformer, which is derived by unrolling a stabilized variant of the expectation-maximization Gerchberg-Saxton (EM-GS) algorithm. Specifically, each layer of the proposed URformer incorporates three trainable modules: 1) a learnable filter implemented by a neural network that replaces the fixed Bessel function ratio in the classic EM-GS algorithm; 2) a trainable gating mechanism that adaptively combines classic and model-based updates to ensure training stability; and 3) a efficient channel Transformer block that learns to correct residual errors by capturing non-local dependencies across the channel matrix. Numerical results demonstrate that the proposed URformer significantly outperforms classic iterative algorithms and conventional black-box neural networks with less pilot overhead.
@@ I Introduction
The demand for ever-increasing data rates and connection density in future 6G networks has spurred research into revolutionary technologies that transcend the limitations of classical radio-frequency (RF) systems [1]. Concurrently, quantum information technologies are rapidly maturing, offering powerful new paradigms for sensing, computation, and communication [2]. At the intersection of these fields lies the Rydberg atomic quantum receiver (RAQR) [3], an emerging quantum sensor technology that utilizes the extreme sensitivity of highly-excited Rydberg atoms to measure electromagnetic fields with quantum-limited precision [4]. Unlike classical RF receivers, which rely on antennas and electronic components that inevitably introduce thermal noise, RAQRs employ an all-optical readout mechanism, such as electromagnetically induced transparency (EIT), to detect changes in atomic quantum states induced by an incident electromagnetic field, making them virtually immune to thermal noise. This combined with their vast frequency tunability from direct-current to Terahertz [3], and their direct traceability to the international system of units without requiring calibration [5], makes them a compelling candidate for next-generation wireless systems.
Integrating RAQRs into wireless communication systems introduces a fundamental modeling departure from classical RF systems, as their signal processing is governed by a non-linear model where the measurement is proportional to the magnitude of the combined electromagnetic field. This magnitude-only measurement creates a challenging phase retrieval problem, making the task of channel acquisition especially difficult. Foundational solutions to this phase retrieval problem are based on classic iterative algorithms, e.g., the Gerchberg-Saxton (GS) and its expectation-maximization (EM) variants [4]. Despite their principled design, these methods suffer from two major drawbacks. First, their fixed mathematical structure is rigid and cannot adapt to non-ideal characteristics of a physical RAQR system. Second, they function as local search methods within a highly non-convex optimization landscape, often getting trapped in suboptimal local minima.
@@ II System Model and Problem Formulation
Figure 1: Uplink multi-user systems with Rydberg atomic quantum receivers.
As illustrated in Fig. 1, we consider an uplink multi-user multiple-input-multiple-output (MIMO) system where $K$ single-antenna users transmit signals to an RAQR equipped with a uniform linear array (ULA) of $M$ Rydberg atomic antennas with inter-antenna spacing $d$. Following the clustered Saleh-Valenzuela (S-V) model, the channel vector $\mathbf{h}_{k}\in\mathbb{C}^{M\times 1}$ from the $k$-th user to the RAQR is formulated as a sum over $L$ clusters with each cluster containing $C_{l}$ rays, which can be expressed as
@@
where $N_{\text{ray}}$ is the total number of rays, $\beta_{k}$ denotes the large-scale fading coefficient for the $k$-th user, $g_{l,c}^{(k)}$ and $\theta_{l,c}^{(k)}$ are the complex gain and direction of arrival (DoA) of the $c$-th ray in the $l$-th cluster for user $k$, respectively. The array steering vector $\mathbf{a}(\theta_{l,c}^{(k)})$ for each ray is given by
@@
Inspired by the measurement model of quantum physics, the RAQR operates based on a four-level atomic energy scheme ($|1\rangle,|2\rangle,|3\rangle,|4\rangle$) and employs an all-optical interferometric or holographic measurement technique [4]. In this scheme, a known local oscillator (LO) RF field interferes with the unknown signal field from the users within the atomic vapor cell. The total RF field at the $m$-th antenna is the superposition of the two, which can be expressed as
$E_{\text{total},m}=[\mathbf{y}_{p}]_{m}+[\mathbf{b}]_{m},$ (4)
where $\mathbf{b}\in\mathbb{C}^{M\times 1}$ denotes the received reference signal of the LO field at the receiver array.
The core measurement mechanism of this architecture is the optical readout of the Autler-Townes (AT) splitting of the EIT spectrum [3, 7]. The observed frequency splitting $\Delta f_{m}$ at the $m$-th antenna by employing photodetector array (PDA) is directly proportional to the total Rabi frequency $\Omega_{R,m}$:
@@
where $\lambda_{c}$ and $\lambda_{p}$ are the wavelengths of the coupling and probe lasers, respectively. The Rabi frequency $\Omega_{R,m}$ is directly proportional to the magnitude of the total electric field acting on the atoms, which is given by [4]
@@
where $\mu_{34}$ is the electric dipole moment of the $|3\rangle\leftrightarrow|4\rangle$ Rydberg transition, a key quantum parameter quantifying the atom-field coupling strength. Combining (5) and (6), we establish that the final measurement $z_{m,p}$ at the photodetector is proportional to the magnitude of the total electric field:
$z_{m,p}\propto\Delta f_{m}\propto\Omega_{R,m}\propto|[\mathbf{H}\mathbf{s}_{p}]_{m}+[\mathbf{b}]_{m}|,$ (7)
For modeling simplicity, all system-dependent proportionality constants are normalized to unity. Including an additive noise term $\mathbf{w}_{p}$, the measurement vector across all $M$ antennas at pilot slot $p$ is given by [^1] Note that a more complex RAQR architecture known as the superheterodyne receiver can yield a different signal model [3, 8]. Under the condition of a strong LO field and a weak signal field, the atomic response can be linearized via a Taylor expansion around the operating point of LO [8]. This results in an equivalent linear baseband model $\mathbf{v}=\mathbf{\Theta}\mathbf{H}\mathbf{s}+\mathbf{w}$, where $\mathbf{\Theta}$ is a diagonal matrix containing the quantum transduction gains and phases. In this work, we focus on the more fundamental and challenging non-linear phase retrieval problem arising from direct AT-splitting measurement schemes.
$\mathbf{z}_{p}=|\mathbf{H}\mathbf{s}_{p}+\mathbf{b}+\mathbf{w}_{p}|,$ (8)
where the noise vector $\mathbf{w}_{p}$ is modeled as $\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$. For an RAQR, this noise is dominated by intrinsic sources rather than thermal noise from environments, which is a combination of blackbody radiation noise (BRN), denotes photon shot noise (PSN), and electronic thermal noise (ETN). This noise components constitutes the standard quantum limit and is typically orders of magnitude lower than the thermal noise floor of classical receivers [4, 9].
The objective is to estimate the channel matrix $\mathbf{H}\in\mathbb{C}^{M\times K}$ from the sequence of magnitude measurement vectors $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{P}]\in\mathbb{R}^{M\times P}$, given the known pilot matrix $\mathbf{s}_{p}$ and the LO channel vector $\mathbf{b}$. This constitutes a structured MIMO biased phase retrieval problem, which can be formulated as
$\hat{\mathbf{H}}=\arg\min_{\mathbf{H}\in\mathbb{C}^{M\times K}}\frac{1}{MP}\sum_{p=1}^{P}\|\mathbf{z}_{p}-|\mathbf{H}\mathbf{s}_{p}+\mathbf{b}|\|_{2}^{2}.$ (9)
A key implication of the non-linear measurement model in Eq. (9) is the breakdown of traditional pilot-based channel estimation strategies. In classical linear MIMO systems, the use of orthogonal pilot sequences allows the receiver to decouple the multi-user channel [10]. This principle fails in RAQRs, where the non-linear magnitude operation $|\cdot|$ is applied to the superposition of all signals of users. Due to the property that $|\sum_{k}\mathbf{h}_{k}s_{k,p}|\neq\sum_{k}|\mathbf{h}_{k}s_{k,p}|$, the orthogonality of the pilot sequences is lost after the measurement is taken. Consequently, the measurement vector for any given pilot transmission is an inseparable function of all channels of users. This invalidates the user-by-user estimation approach and necessitates a joint estimation of the entire channel $\mathbf{H}$.
@@ III-A Baseline Phase Retrieval Algorithms
The conventional approach to solving the biased phase retrieval problem in Eq. (8) relies on iterative algorithms [4]. Typically, the GS algorithm starts with an initial channel guess $\hat{\mathbf{H}}^{(0)}$, and iteratively refines it by alternating between estimating the phase of signal from the current channel estimate. A statistically robust extension is the EM-GS algorithm, which is to weight the reconstructed signal not just with the estimated phase, but also by a factor that depends on the local signal-to-noise ratio (SNR) proxy $\boldsymbol{\kappa}$. This weighting is performed by the ratio of modified Bessel functions $R(\boldsymbol{\kappa})$, which acts as a soft filter to suppress noisy components. The update rule in the $t$-th iteration is given by [4]
$\mathbf{Y}_{\text{complex}}^{(t-1)}=\hat{\mathbf{H}}^{(t-1)}\mathbf{S}^{T}+\mathbf{B},$ (10)
$\mathbf{\boldsymbol{\kappa}}^{(t-1)}=\frac{2\mathbf{Z}\circ|\mathbf{Y}_{\text{complex}}^{(t-1)}|}{\sigma^{2}},$ (11)
$\mathbf{Y}_{\text{rec, EM-GS}}^{(t-1)}=\mathbf{Z}\circ e^{i\angle\mathbf{Y}_{\text{complex}}^{(t-1)}}\circ R(\mathbf{\boldsymbol{\kappa}}^{(t-1)}),$ (12)
$\hat{\mathbf{H}}^{(t)}=(\mathbf{Y}_{\text{rec, EM-GS}}^{(t-1)}-\mathbf{B})(\mathbf{S}^{T})^{\dagger}.$ (13)
The performance of the EM-GS algorithm is fundamentally limited by its core filtering function $R(\boldsymbol{\kappa})$. This fixed filter, derived under an ideal Gaussian noise model, is non-adaptive to practical system impairments and prone to error propagation in low-SNR conditions. Moreover, as a local search method navigating a highly non-convex optimization landscape, the algorithm is susceptible to getting trapped in local optima, making its performance critically dependent on the initial estimate. These inherent limitations motivate our approach to replace the rigid components of the algorithm with learnable modules within a deep unrolling framework.
@@ III-B Proposed URformer Architecture
Figure 2: Proposed URformer architecture for channel estimation.
As illustrate in Fig. 2, we propose a URformer architecture to solve this non-linear estimation problem by unrolling a $T_{\text{GS}}$-iteration EM-GS algorithm into a $T_{\text{UR}}$-layer neural network, where the fixed computational blocks in EM-GS algorithm, e.g., $R(\boldsymbol{\kappa})$, are replaced by learnable modules. Note that the number of unrolled layers $T_{\text{UR}}$, is typically much smaller than the iterations required by the classic algorithm, stemming from the data-driven learnable modules. Each URformer layer performs one iteration of a modified and enhanced EM-GS algorithm. A critical challenge in unrolling this algorithm is its potential for instability during training. To address this, we introduce a trainable gating mechanism that allows the network to learn a stable update policy. Each URformer layer takes the channel estimate $\hat{\mathbf{H}}^{(t-1)},(1\leq t\leq T_{\text{UR}})$ of the previous layer and produces an updated estimate $\hat{\mathbf{H}}^{(t)}$. A single URformer layer consists of the following three modules.
@@ III-B1 Gated Filtering Module
This module replaces the fixed Bessel function filter with a learnable FilterNet and adds a stabilization gate. First, the complex signal is estimated as
$\mathbf{Y}_{\text{complex}}=\hat{\mathbf{H}}^{(t-1)}\mathbf{S}^{T}+\mathbf{B}.$ (14)
The phase is extracted as $\mathbf{\Phi}=\angle\mathbf{Y}_{\text{complex}}$. A direct GS-style reconstruction is given by
$\mathbf{Y}_{\text{direct}}=\mathbf{Z}\circ e^{i\boldsymbol{\Phi}}.$ (15)
Simultaneously, the parameter $\mathbf{K}$ is computed as in (11). We employ a small MLP, termed $\text{FilterNet}_{\psi}$ with trainable weights $\psi$}, to learn a filtering function:
$R_{\text{learned}}=\text{FilterNet}_{\psi}(\mathbf{K}).$ (16)
The filtered reconstruction is given by
$\mathbf{Y}_{\text{filtered}}=\mathbf{Y}_{\text{direct}}\circ R_{\text{learned}}.$ (17)
A learnable gating parameter $\alpha^{(t)}$ combines the two reconstructions, which can be expressed as
$\mathbf{Y}_{\text{rec}}=\alpha^{(t)}\mathbf{Y}_{\text{filtered}}+(1-\alpha^{(t)})\mathbf{Y}_{\text{direct}},$ (18)
where $\alpha^{(t)}=\sigma(g^{(t)})$ denotes the sigmoid activation of a trainable parameter $g^{(t)}$.

### v2

@@ Abstract
A model-driven deep learning framework is proposed for channel estimation in Rydberg atomic quantum receivers (RAQRs) based on the measurement of holographic snapshots. Specifically, we develop a Transformer-based unrolling architecture, termed URformer, to solve the non-linear biased phase retrieval problem, which is derived by unrolling a stabilized variant of the expectation-maximization Gerchberg-Saxton (EM-GS) algorithm. Each layer of the proposed URformer incorporates three trainable modules: 1) a learnable filter network that replaces the fixed Bessel kernel in the classic EM-GS algorithm; 2) a trainable gating mechanism that adaptively combines classic updates to ensure training stability; and 3) an efficient channel Transformer module that learns to correct residual errors by capturing non-local channel dependencies. Numerical results demonstrate that the proposed URformer significantly outperforms classic iterative algorithms and conventional black-box neural networks with less pilot overhead.
@@ I Introduction
The demand for ever-increasing data rates and connection density in future 6G networks has spurred research into revolutionary antenna technologies that transcend the limitations of classical radio-frequency (RF) systems [1, 2]. Concurrently, quantum information technologies are rapidly maturing, offering new paradigms for sensing, computation, and communication [3]. At the intersection of these fields lie Rydberg atomic quantum receivers (RAQRs) [4], an emerging quantum sensor technology that utilizes the extreme sensitivity of highly-excited Rydberg atoms to measure electromagnetic fields with quantum-limited precision [5].
Integrating RAQRs into wireless communication systems necessitates a distinct detection architecture that departs from classical RF systems. Specifically, the superheterodyne receiver is the popular architecture in RAQRs [6], which can exhibit an equivalent linear baseband signal model via a Taylor expansion around the operating point of the strong local oscillator (LO) [4, 7]. However, the superheterodyne RAQRs impose significant hardware complexity for large-scale arrays, as they necessitate dedicated high-bandwidth photodetectors and synchronous demodulation circuitry for each array element to resolve the time-varying intermediate frequency signal. Consequently, in this work, we develop the holographic snapshot-based RAQR architecture to enable scalable parallel readout in RAQRs, which is achieved by utilizing an optical imaging array to capture the spatial intensity profile of the probe laser beam [8], effectively allowing atomic antenna elements to be read out simultaneously in a single exposure. This RAQR architecture also introduces a new signal processing challenge where the magnitude-only measurement is obtained in digital domain and hence results in a non-linear phase retrieval problem, making the task of channel acquisition especially difficult. Foundational solutions to this phase retrieval problem are based on classic iterative algorithms, e.g., the Gerchberg-Saxton (GS) and its expectation-maximization (EM) variants [5]. Despite their principled design, these methods suffer from two major drawbacks. First, their fixed mathematical structure is rigid and cannot adapt to non-ideal characteristics of a physical RAQR system. Second, they function as local search methods within a highly non-convex optimization landscape, often getting trapped in suboptimal local minima.
@@ II System Model and Problem Formulation
Figure 1: Uplink multi-user systems with the RAQR. By capturing the intensity profile of the probe laser with an optical field imaging array, i.e., holographic snapshots, all spatial channels are detected simultaneously, which reduces the hardware complexity compared to the superheterodyne RAQR.
As illustrated in Fig. 1, we consider an uplink multi-user multiple-input-multiple-output (MIMO) system where $K$ single-antenna users transmit signals to an RAQR equipped with a uniform linear array (ULA) of $M$ Rydberg atomic antennas with inter-antenna spacing $d$.
II-1 Channel Model
Considering the clustered Saleh-Valenzuela model [10], the channel vector $\mathbf{h}_{k}\in\mathbb{C}^{M\times 1}$ from the $k$-th user to the RAQR is formulated as a sum over $L$ clusters with each cluster containing $C_{l}$ rays, which can be expressed as
@@ II-1 Channel Model
where $N_{\text{ray}}$ is the total number of rays. $g_{l,c}^{(k)}$ and $\theta_{l,c}^{(k)}$ are the complex path gain and direction of arrival (DoA) of the $c$-th ray in the $l$-th cluster for user $k$, respectively. The array steering vector $\mathbf{a}(\theta_{l,c}^{(k)})$ for each ray is given by
@@
II-2 Holographic Snapshot Measurement Model
The RAQR operates based on a four-level atomic energy scheme ($|1\rangle\to|2\rangle\to|3\rangle\leftrightarrow|4\rangle$) [8], [11]. A reference LO field vector $\mathbf{b}\in\mathbb{C}^{M\times 1}$ interferes with the impinging multi-user signal field $\mathbf{y}_{p}$ within the vapor cell. The total RF electric field vector across the array is the coherent superposition:
$\mathbf{E}_{\text{total}}=\mathbf{y}_{p}+\mathbf{b}.$ (4)
To achieve scalable parallel readout, we employ an optical field imaging array, e.g., a high-resolution CMOS or CCD camera, to capture the spatial intensity profile of the probe laser beam. By operating on the linear slope of the electromagnetically induced transparency (EIT) resonance, the imaging array performs a frequency-to-intensity mapping, where the local Autler-Townes (AT) frequency splitting $\Delta f$ is linearly translated into an intensity variation recorded by the sensor pixels. Specifically, the observed frequency splitting $\Delta f_{m}$ at the $m$-th pixel antenna is given by
@@ II-2 Holographic Snapshot Measurement Model
where $\lambda_{c}$ and $\lambda_{p}$ are the wavelengths of the coupling and probe lasers, respectively. $\Omega_{R,m}$ denotes the Rabi frequency which is directly linked to the magnitude of the total electric field acting on the atoms:
@@
where $\mu_{34}$ is the electric dipole moment of the Rydberg transition and $\hbar$ is the reduced Planck constant.
We define the measurement vector $\mathbf{z}_{p}$ at pilot slot $p$ as a holographic snapshot, representing the static magnitude of the spatial interference pattern. Combining Eqs. (4)-(6), the array-level measurement model is formulated as
$\mathbf{z}_{p}=\frac{\lambda_{c}\mu_{34}}{2\pi\hbar\lambda_{p}}\left|\mathbf{H}\mathbf{s}_{p}+\mathbf{b}+\mathbf{w}_{p}\right|,$ (7)
where the noise vector $\mathbf{w}_{p}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$ accounts for the intrinsic system noise, which is dominated by photon shot noise and atomic blackbody radiation, along with the sensor readout noise inherent to the imaging array. Let $\mathcal{G}$ denote the atomic transduction gain $\mathcal{G}=\frac{\lambda_{c}\mu_{34}}{2\pi\hbar\lambda_{p}}$.
II-3 Problem Formulation
The objective is to estimate the channel matrix $\mathbf{H}\in\mathbb{C}^{M\times K}$ from the sequence of magnitude-only holographic snapshots $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{P}]\in\mathbb{R}^{M\times P}$. Given the known pilot matrix $\mathbf{S}$ and the LO field $\mathbf{b}$, this constitutes a structured MIMO biased phase retrieval problem:
$\hat{\mathbf{H}}=\arg\min_{\mathbf{H}\in\mathbb{C}^{M\times K}}\frac{1}{MP}\sum_{p=1}^{P}\left\|\mathbf{z}_{p}-\mathcal{G}\left|\mathbf{H}\mathbf{s}_{p}+\mathbf{b}\right|\right\|_{2}^{2}.$ (8)
A key implication of the non-linear measurement model in Eq. (8) is the breakdown of traditional pilot-based channel estimation strategies. In classical linear MIMO systems, the use of orthogonal pilot sequences allows the receiver to decouple the multi-user channel. This principle fails in RAQRs, where the non-linear magnitude operation $|\cdot|$ is applied to the superposition of all signals of users. Due to the property that $|\sum_{k}\mathbf{h}_{k}s_{k,p}|\neq\sum_{k}|\mathbf{h}_{k}s_{k,p}|$, the orthogonality of the pilot sequences is lost after the measurement is taken. Consequently, the measurement vector for any given pilot transmission is an inseparable function of all channels of users. This invalidates the user-by-user estimation approach and necessitates a joint estimation of the entire channel $\mathbf{H}$.
@@ III-A Baseline Phase Retrieval Algorithms
The conventional approach to solving the biased phase retrieval problem in Eq. (8) relies on iterative algorithms [5]. Typically, the GS algorithm starts with an initial channel guess $\hat{\mathbf{H}}^{(0)}$, and iteratively refines it by alternating between estimating the phase of signal from the current channel estimate. A statistically robust extension is the EM-GS algorithm, which is to weight the reconstructed signal not just with the estimated phase, but also by a factor that depends on the local signal-to-noise ratio (SNR) proxy $\boldsymbol{\kappa}$. This weighting is performed by the ratio of modified Bessel functions $R(\boldsymbol{\kappa})$, which acts as a soft filter to suppress noisy components. Given the known LO reference field matrix $\mathbf{B}\in\mathbb{C}^{M\times P}$ by stacking the LO vector $\mathbf{b}$ across $P$ pilot time slots, the update rule in the $t$-th iteration is given by [5]
$\mathbf{Y}^{(t-1)}=\hat{\mathbf{H}}^{(t-1)}\mathbf{S}^{T}+{\mathbf{B}},$ (9)
$\mathbf{\boldsymbol{\kappa}}^{(t-1)}=\frac{2\mathbf{Z}{\cdot}|\mathbf{Y}^{(t-1)}|}{\sigma^{2}},$ (10)
$\mathbf{Y}_{\text{EM-GS}}^{(t-1)}=\mathbf{Z}{\cdot}e^{i\angle\mathbf{Y}^{(t-1)}}{\cdot}R(\mathbf{\boldsymbol{\kappa}}^{(t-1)}),$ (11)
$\hat{\mathbf{H}}^{(t)}=(\mathbf{Y}_{\text{EM-GS}}^{(t-1)}-\mathbf{B})(\mathbf{S}^{T})^{\dagger}.$ (12)
The performance of the EM-GS algorithm is fundamentally limited by its core filtering function $R(\boldsymbol{\kappa})$. This fixed filter derived under an ideal Gaussian noise model is prone to error propagation in low-SNR conditions. Moreover, as a local search method navigating a highly non-convex optimization landscape, the EM-GS algorithm is susceptible to getting trapped in local optima, making its performance critically dependent on the initial estimate.
@@ III-B Proposed URformer Architecture
As illustrated in Fig. 2, we propose a URformer architecture to solve the non-linear channel estimation problem by unrolling a $T_{\text{GS}}$-iteration EM-GS algorithm into a $T_{\text{UR}}$-layer neural network, where the fixed computational blocks in EM-GS algorithm, e.g., $R(\boldsymbol{\kappa})$, are replaced by learnable modules. Note that the number of unrolled layers $T_{\text{UR}}$ is typically much smaller than the $T_{\text{GS}}$ iterations required by the classic algorithm, stemming from the data-driven learnable modules. A critical challenge in unrolling the EM-GS algorithm is its potential for instability during training. To address this, we introduce a trainable gating mechanism that allows the network to learn a stable update policy. Each URformer layer takes the channel estimate $\hat{\mathbf{H}}^{(t-1)}(1\leq t\leq T_{\text{UR}})$ of the previous layer and produces an updated estimate $\hat{\mathbf{H}}^{(t)}$. The initial channel estimate $\hat{\mathbf{H}}^{(0)}$ is generated as a random complex Gaussian matrix with zero mean and normalized variance. A single URformer layer consists of the following three modules.
@@ III-B1 Gated Filtering Module
This module replaces the fixed Bessel function filter with a learnable FilterNet and adds a stabilization gate. First, the signal $\mathbf{Y}^{(t-1)}$ is estimated according to Eq. (9), and then the phase is extracted as $\mathbf{\Phi}^{(t-1)}=\angle\mathbf{Y}^{(t-1)}$. We carry out a direct GS-style reconstruction as
$\mathbf{Y}_{\text{direct}}^{(t-1)}=\mathbf{Z}{\cdot}e^{i\boldsymbol{\Phi}^{(t-1)}}.$ (13)
Simultaneously, the parameter $\mathbf{\boldsymbol{\kappa}}^{(t-1)}$ is computed as in Eq. (10). We employ a $\text{FilterNet}_{{\psi}^{(t-1)}}$ with trainable weights ${\psi}^{(t-1)}$ as an element-wise operator to learn a filtering function:
$R_{\text{learned}}^{(t-1)}=\text{FilterNet}_{{\psi}^{(t-1)}}(\mathbf{\boldsymbol{\kappa}}^{(t-1)}),$ (14)
where the FilterNet is designed as a compact two-layer multi-layer perceptron (MLP) that maps the scalar SNR proxy $\kappa^{(t-1)}$ to a filtering coefficient in the range $(0,1)$.
Then, the filtered reconstruction is given by
$\mathbf{Y}_{\text{filtered}}^{(t-1)}=\mathbf{Y}_{\text{direct}}^{(t-1)}{\cdot}R_{\text{learned}}^{(t-1)}.$ (15)
A learnable gating parameter $\alpha^{(t-1)}$ combines the two reconstructions, which can be expressed as
$\mathbf{Y}_{\text{rec}}^{(t-1)}=\alpha^{(t-1)}\mathbf{Y}_{\text{filtered}}^{(t-1)}+(1-\alpha^{(t-1)})\mathbf{Y}_{\text{direct}}^{(t-1)},$ (16)
where $\alpha^{(t-1)}=\sigma(g^{(t-1)})$ denotes the sigmoid activation of a trainable parameter $g^{(t)}$.